# Protect Before Generate: Error Correcting Codes within Discrete Deep Generative Models

### v1

@@ 4 Coded DVAE
This section extends the previously described DVAE, introducing an ECC over $\boldsymbol{m}$. We refer to this model as coded DVAE. In ECCs, we augment the dimensionality of the binary latent space from $M$ to $D$ in a controlled and deterministic manner, where $R=M/D$ is the coding rate. An ECC is typically designed so that the $2^{M}$ possible codewords are separated as much as possible in the space of binary vectors of $D$ bits. This facilitates algorithms in detecting and/or correcting errors by searching for the nearest code word. A random choice of the codewords brings what is known as a random block code [25]. While they are known to be very robust and amenable to theoretical analysis, their lack of structure makes them computationally intractable since we have to rely on codeword enumeration during the encoding/decoding process. In Appendix J, we include the formulation of a random code’s encoding/decoding process within the DVAE model.
@@
Instead, we adopt a much simpler linear coding scheme, namely repetition codes. In a repetition code, each bit of the original message $\boldsymbol{m}$ is repeated multiple times to create the encoded message $\boldsymbol{c}$. Intuitively, the more times an information bit is repeated, the better it is protected. Our experiments consider uniform $(M,D)$ repetition codes where all bits are repeated $L$ times, resulting in codewords of dimension $D=ML$ and a coding rate of $R=1/L$. Note that repetition codes represent a special case of linear ECCs since each codeword can be deterministically computed by multiplying a binary vector $\boldsymbol{m}$ by an $M\times D$ generator matrix $\mathbf{G}$, such that $\boldsymbol{c}=\boldsymbol{m}^{T}\mathbf{G}$, where $u$-th row, with $u=1,\ldots,M$, has entries equal to one at columns $L(u-1)+1,L(u-1)+2,\dots,Lu$, and zero elsewhere. For example, for $M=3$ and $L=2$, the generator matrix of the $(3,6)$ repetition code is
@@
The generative process of the coded DVAE follows similarly to the uncoded case, and it is represented in Fig. 1(d). We assume the same prior distribution $p(\boldsymbol{m})$, but in this case the samples $\boldsymbol{m}$ are deterministically encoded using $\mathbf{G}$. Now, the smoothing $\boldsymbol{z}$ transformations are defined over $\boldsymbol{c}$
$p(\boldsymbol{z}|\boldsymbol{c})=\prod_{j=1}^{D}p(z_{j}|c_{j}),\quad p(z_{j}|c_{j})$ $=\frac{e^{-\beta(z_{j}-c_{j})}}{Z_{\beta}},$ (7)
@@
for $z_{j}\in[0,1]$, $c_{j}\in\{0,1\}$ and $Z_{\beta}=(1-e^{-\beta})/\beta$. The likelihood $p(\boldsymbol{x}|\boldsymbol{z})$ is again of the form $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$. Note that, compared to the uncoded case, we have a larger input dimensionality to the decoder NN $f_{\boldsymbol{\theta}}(\boldsymbol{z})$. When comparing uncoded vs. coded DVAEs, the structure of the decoder NN $f_{\boldsymbol{\theta}}(\boldsymbol{z})$ (detailed in Appendix A) is equal in both cases except for the first Multilayer Perceptron (MLP) layer that attacks the input $\boldsymbol{z}$. Therefore, if a rate $R=1/L$ repetition code is used, the number of additional parameters of the $f_{\boldsymbol{\theta}}(\boldsymbol{z})$ NN is given by $(L-1)\times h$, where $h$ is the dimension of the first hidden space of $f_{\boldsymbol{\theta}}(\boldsymbol{z})$.
Variational family and inference
The repetition code introduces correlations between the bits in $\boldsymbol{c}$ that we will exploit to obtain an improved variational bound. We again assume a variational family factorizing as
$q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{c})$ (8)
@@
where $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})=\prod_{u=1}^{M}q_{\boldsymbol{\eta}}(m_{u}|\boldsymbol{x})$ is computed in two steps. First, we construct an encoder NN $\mathbf{g}_{\boldsymbol{\eta}}(\boldsymbol{x})$ similar to that of (2), that retrieves the probabilities of the bits in $\boldsymbol{c}$ from $\boldsymbol{x}$ without exploiting the correlations introduced by the repetition code:
$q^{u}_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$ $=\prod_{j=1}^{D}\text{Ber}(g_{j,\boldsymbol{\eta}}(\boldsymbol{x})),$ (9)
where the $u$ superscript serves as a reminder that this posterior does not exploit the redundancy introduced by the ECC.
@@
Now, we utilize the known redundancy introduced by the ECC to constrain the solution of $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x})$, given that each bit from $\boldsymbol{m}$ has been repeated $L$ times to create $\boldsymbol{c}$. To do so, we follow a soft decoding approach, where the marginal posteriors of the information bits are derived from the marginal posteriors of the encoded bits, exploiting the repetition code’s known structure. In the case of repetition codes, we compute the all-are-zero and the all-are-ones products of probabilities of the bits in $\boldsymbol{c}$ that are copies of the same message bit and renormalize as
$q(m_{u}=1|\boldsymbol{x})$ $=\frac{1}{Z}\prod_{j=L(u-1)+1}^{Lu}g_{j,\boldsymbol{\eta}}(\boldsymbol{x})\doteq\frac{g^{+}_{u,\boldsymbol{\eta}}(\boldsymbol{x})}{Z},$ (10)
$q(m_{u}=0|\boldsymbol{x})$ $=\frac{1}{Z}\prod_{j=L(u-1)+1}^{Lu}(1-g_{j,\boldsymbol{\eta}}(\boldsymbol{x}))\doteq\frac{g^{-}_{u,\boldsymbol{\eta}}(\boldsymbol{x})}{Z},$ (11)
for $u=1,\ldots,M$ and $Z=\left(g^{+}_{u,\boldsymbol{\eta}}(\boldsymbol{x})+g^{-}_{u,\boldsymbol{\eta}}(\boldsymbol{x})\right)^{-1}$. This approach can be seen as a soft majority voting strategy, enabling the recovery of the original information vector even if some bits in the inferred encoded word are corrupted. All operations in (10) preserve the gradients concerning the parameters in the encoder $\mathbf{g}_{\boldsymbol{\eta}}(\boldsymbol{x})$. We implement them in the log domain for stability.
@@
When compared to the uncoded case, as in the likelihood term, we consider the same NN structure for the encoder $\mathbf{g}_{\boldsymbol{\eta}}(\boldsymbol{x})$ where both cases only differ in the last MLP layer. The additional overhead in the coded cases requires $(L-1)\times h^{\prime}$ parameters in the last layer, where $h^{\prime}$ is the dimension at the output of the last layer. In Appendix H.2, we conduct an ablation study on the number of trainable parameters to demonstrate that the improvement in performance does not stem from this increase in the number of parameters, but rather from the incorporation of the ECC in the latent space.
Algorithm 2 Training the coded DVAE with repetition codes.
1:  Input: training data $\boldsymbol{x}_{i}$, matrix $\mathbf{G}$.
2:  repeat
6:     $\boldsymbol{z}\leftarrow$ sample from (5)
8:     Compute ELBO according to (3)
9:     $\boldsymbol{\theta},\boldsymbol{\eta}\leftarrow Update(ELBO)$
10:  until convergence
Soft encoding for efficient reparameterization
Given the variational family in (8), the ELBO matches the expression in (3). However, the reparameterization trick in (5) requires independent bits, which is not the case in $\boldsymbol{c}$. To efficiently circumvent this issue during training, we employ a soft encoding approach. With soft encoding, a marginal probability is computed for each bit in the codeword $\boldsymbol{c}$, taking into account the structure of the ECC and the marginal probabilities of the information bits. For a repetition code, this involves simply replicating the posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})=\prod_{u=1}^{M}q_{\boldsymbol{\eta}}(m_{u}|\boldsymbol{x})$ for each copy of the same information bit. Hence, we treat the bits in $\boldsymbol{c}$ as independent but distributed according to $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. The Algorithm 2 shows the training pseudo-code.
When marginalizing $\boldsymbol{c}$ using the soft encoding marginals, we disregard potential correlations between the coded bits. For instance, with repetition codes, sampling from the marginals could produce inconsistent bits, leading to an invalid codeword. However, since we do not sample the coded bits during training but instead propagate their marginal probabilities, we consider this approximation to have minimal negative impact. In fact, it can be seen as a form of probabilistic dropout, which enhances robustness during training. It is important to note that when sampling from the generative model in test time, we use hard bits encoded into valid codewords, yielding visually appealing samples, indicating that our training approach is reliable.
Related work
While the use of deep neural networks and generative models in digital communications problems have been profusely reported in recent years (see [36], [5], [9], [33], and [26] for representative examples), the use of ECC techniques as a design tool in machine learning is scarce. The most prominent example is [7], where the authors address multiclass learning problems via ECCs. In [1], the authors proposed using linear codes for applications that may require removing the trace of a sample from the system, e.g., a user requests their data to be deleted or corrupted data is discovered. They address a regression problem by introducing a coded learning protocol that employs linear encoders to divide the training data before the learning phase. More recently, in [35] the authors introduced ECCs to improve code-to-code translation using transformers.
4.1 Reconstruction
We first evaluate the model’s performance of reconstructing data by examining its uncoded and coded versions across different configurations, varying the number of information bits and code rates. The introduction of the repetition code led to improved reconstruction and smaller KL values, indicating that the posterior latent features are disentangled and less correlated. In Appendices B, C, D and E, we show the behavior of the ELBO loss function for all the models and datasets.
Image reconstruction quality.   In the table included in Fig. 2, we first quantify the quality of the reconstructions in FMNIST by measuring the Peak Signal-To-Noise Ratio (PSNR) in the test set. The results for the rest of the datasets are provided in sections B, C, D and E of the Appendix. In all the cases, the coded models yield higher PSNR values than their uncoded counterparts, indicating a superior performance in reconstruction. This improvement is also evident by visual inspection of Fig. 2, where the coded models exhibit a greater ability to capture details in the images for the same latent dimension. We observe a general improvement in PSNR as we increase the number of information bits, i.e., as we augment the latent dimensionality of the model. This increase in the number of available latent vectors provides greater flexibility, enabling the models to capture the underlying structure of the data more effectively. We also observe a general improvement in PSNR as we decrease the code rate, i.e., as we add more redundancy. Note that adding redundancy does not increase the model’s flexibility, since the information bits determine the number of latent vectors. However, coded models yield more accurate and detailed reconstructions.
@@ 4.1 Reconstruction
Figure 2: Reconstruction performance over the test set in FMNIST. The figure at the left shows an example of reconstructed test images obtained with different model configurations. Observe that more details are visualized as we increase the number of bits in the latent space and decrease the coding rate. The table at the right includes reconstruction metrics. Acc is the semantic accuracy and Conf. Acc the confident semantic accuracy. Entropy is the average entropy of $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ in the test set.
Semantic accuracy.   As the PSNR operates at the pixel level, it does not account for the semantic errors committed by the model. For example, if the model incorrectly reconstructs a nine instead of a four in the MNIST dataset, the PSNR may still yield a large value due to the similarity between the two images. Nonetheless, this would represent a severe failure in correct reconstruction of the intended class. Therefore, we additionally evaluate the reconstruction accuracy, ensuring that the model successfully reconstructs images within the same class as the original ones. For this purpose, we trained an image classifier for each dataset and compared the reconstructed images’ predicted labels against the originals’ ground truth labels. Additionally, we provide a confident reconstruction accuracy. While the reconstruction accuracy is computed across the entire dataset partitions, for the confident accuracy, we only consider those images projected into a latent vector with a probability exceeding 0.4.111Namely, we do not count errors when the Maximum a Posteriori (MAP) value of $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ is below 0.4.. Results for FMNIST are detailed in the table included in Fig.2, and corresponding results for MNIST are available in Table 2 within Appendix C. In light of the results, we can conclude that introducing an ECC in the model allows for latent spaces that better capture the semantics of the images while employing the same number of latent vectors, significantly outperforming the uncoded models in terms of accuracy in all the cases.
Figure 3: Example of erroneous reconstructions in FMNIST using the 4 most-probable a posteriori latent vectors. The first column in each image shows the original input to the model, while the second column displays the reconstructions. The $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ probability is indicated in each row.
Posterior uncertainty calibration.   Finally, also in the table included in Fig.2, we report the average entropy of the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ over the test set. The low entropy observed in the uncoded models suggests a low uncertainty when the model projects data points into the latent space, which could be advantageous if the model consistently assigned high probability to the correct latent vectors. However, the semantic accuracy results demonstrate this is not true in the uncoded model. In other words, the uncoded variational family projects images into the wrong class with high confidence. This indicates the uncertainty of the uncoded case is severely miscalibrated.
Coded models, on the other hand, improve semantic accuracy and present a larger entropy. This suggests that i) the coded DVAE is aware that multiple latent vectors might be related to the image class and ii) that the model posterior shows large uncertainties (high entropy) for certain images for which the model has not properly identified the class. We illustrate this in Fig. 3, where we show some images that were selected so that the MAP latent word from $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ induces class reconstruction errors. We display the reconstruction of the 4 most probable latent vectors and their corresponding probabilities. Observe that the uncoded model is confident no matter the reconstruction outcome while, in the coded posterior, the uncertainty is much larger. These results are also observed for MNIST, indicating that the posterior distribution in coded models exhibits a better uncertainty calibration. Note also that the increase in the number of latent bits (from 8 to 10) does not result in an excessive increase in the entropy despite the exponential growth of the number of vectors.
@@ 4.2 Generation
Model BER WER LL train LL test
uncoded 5 0.051 0.195 -266.157 -267.703
coded 5/50 0.011 0.046 -239.379 -241.882
coded 5/80 0.008 0.039 -227.550 -232.992
coded 5/100 0.010 0.049 -238.206 -241.404
uncoded 8 0.089 0.384 -247.964 -249.880
coded 8/80 0.021 0.144 -227.550 -232.992
coded 8/160 0.027 0.189 -228.585 -235.819
coded 8/240 0.037 0.231 -231.679 -238.459
uncoded 10 0.142 0.622 -242.842 -244.997
coded 10/100 0.040 0.321 -222.011 -230.772
coded 10/200 0.044 0.341 -223.748 -234.849
coded 10/300 0.045 0.349 -226.504 -238.647
Figure 4: Evaluation of generation in FMNIST. The figure at the left shows an example of randomly generated, uncurated FMNIST images. The table at the right shows the quantitative results on the evaluation of the BER, WER, and log-likelihood (LL).
@@
In this section, we evaluate the model for the image generation task. In Fig. 4, we show examples of randomly generated images using different model configurations in FMNIST. Results for the rest of the datasets are available in Appendices B, C, D, and E. These results are consistent with the ones obtained in reconstruction since we can observe that the coded models can generate more detailed and diverse images. Both uncoded and coded models generate more intricate and varied images with increased information bits. However, if the number of latent vectors becomes too large for the dataset’s complexity, not all words in the codebook are specialized during model training. This leads to generation artifacts, images where different classes of objects are overlapped. A visual inspection of Fig. 4 suggests these artifacts are more frequent in the uncoded case. Note that, since we are dealing with discrete latent variables, we could simply detect and prune uninformative vectors.
@@
Accuracy metrics in generation.   The improved inference given by the repetition code can also be tested by generating images using the generative model and counting errors using the MAP solution of the variational distribution $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. The table included in Fig. 4 reports the BER and WER for FMNIST. As expected, at the same number of latent bits, the coded models significantly reduce both the BER and WER w.r.t. the uncoded case. Note also that the error rates grow with the number of latent bits, which is expected due to the increased complexity of the inference process. We may commit more errors by taking the MAP, but errors typically fall in consistent reconstructions (latent words that also reconstruct the same type of image), as the results presented in Section 4.1 indicated.
Log-likelihood.   We additionally estimated the log-likelihood (LL). Results for FMNIST with different model configurations, estimated through importance sampling with 300 samples per observation, are presented in the table included in Fig. 4, please refer to Appendix I for further details. We observe that coded models consistently outperform their uncoded counterparts for both train and test sets, aligning with the results previously presented. Thus, we can argue that the introduction of repetition codes in the definition of the model allows for an improved inference and tighter posterior approximation. We draw similar conclusions for other datasets.
We observe a general improvement in LL values as we increase the number of information bits, i.e., as we augment the latent dimensionality of the model and its flexibility. However, reducing the code rate does not lead to an improvement in log-likelihood. We argue that this might indicate overfitting of the decoder, as the LL deteriorates while reconstruction metrics improve. We must note that we use feed-forward networks at the decoder’s input. However, this may not be appropriate for the correlations we present in our coded words. We might overcome this overfitting tendency by using an architecture that properly leverages these coded bits correlations.
Figure 5: Reconstruction results for CIFAR10 (left) and Tiny ImageNet (right).

### v2

@@ 4 Coded-DVAE
This section extends the previously described DVAE, introducing an ECC over $\boldsymbol{m}$. We refer to this model as Coded-DVAE. The corresponding graphical model is shown in Fig. 3, where the diamond node represents a deterministic transformation. In ECCs, we augment the dimensionality of the binary latent space from $M$ to $D$, introducing redundancy in a controlled and deterministic manner, where $R=M/D$ is the coding rate. An ECC is designed to maximize the separation between the $2^{M}$ possible codewords within the $D$-bit binary space, enabling efficient error correction by searching for the nearest valid codeword (see Fig. 1(a)). A random selection of codewords results in random block codes [Shannon, 1948], which are very robust and amenable to theoretical analysis. However, their lack of structure makes them impractical, as encoding and decoding require exhaustive codeword enumeration. Appendix N provides a detailed formulation of the encoding and decoding process for a random code within the Coded-DVAE.
@@
We adopt a much simpler linear coding scheme, namely repetition codes, where each bit of the original message $\boldsymbol{m}$ (i.e., each information bit) is duplicated multiple times to form the encoded message $\boldsymbol{c}$. Intuitively, the more times an information bit is repeated, the better it is protected. Our experiments consider uniform $(M,D)$ repetition codes where all bits are repeated $L$ times, resulting in codewords of dimension $D=ML$ and a coding rate of $R=1/L$. Note that repetition codes represent a special case of linear ECCs since each codeword can be computed by multiplying a binary vector $\boldsymbol{m}$ by an $M\times D$ generator matrix $\mathbf{G}$, such that $\boldsymbol{c}=\boldsymbol{m}^{T}\mathbf{G}$, where $k$-th row, with $k=1,\ldots,M$, has entries equal to one at columns $L(k-1)+1,L(k-1)+2,\dots,Lk$, and zero elsewhere. For example, for $M=3$ and $L=2$, the generator matrix of the $(3,6)$ repetition code is
@@
The generative process of the Coded-DVAE is similar to the uncoded version and is illustrated in Fig. 3. We assume the same prior distribution $p(\boldsymbol{m})$, but in this case the samples $\boldsymbol{m}$ are deterministically encoded using $\mathbf{G}$. Consequently, the smoothing transformations are now defined over $\boldsymbol{c}$
$p(\boldsymbol{z}|\boldsymbol{c})=\prod_{j=1}^{D}p(z_{j}|c_{j}),$ (8)
@@
with $p(z_{j}|c_{j})$ following Eq. (1). The likelihood $p(\boldsymbol{x}|\boldsymbol{z})$ is again of the form $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$. Compared to the uncoded case, the input dimensionality to the decoder $f_{\boldsymbol{\theta}}(\boldsymbol{z})$ is $L$ times larger due to the introduced redundancy. Therefore, the overall architecture of the decoder (detailed in Appendix C) remains unchanged, except for the first Multilayer Perceptron (MLP) layer that processes the input $\boldsymbol{z}$. When using a repetition code with rate $R=1/L$, the additional parameters of $f_{\boldsymbol{\theta}}(\boldsymbol{z})$ amount to $(L-1)\times h$, where $h$ is the dimension of the first hidden layer.
4.1 Variational family and inference
Since $\boldsymbol{c}$ is a deterministic transformation of $\boldsymbol{m}$, its randomness is entirely determined by $\boldsymbol{m}$. Therefore, following the recognition model described in Fig. 3, we assume a variational family factorizing as
$q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{c}),$ (9)
@@ 4.1 Variational family and inference
where $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})=\prod_{j=1}^{M}q_{\boldsymbol{\eta}}(m_{j}|\boldsymbol{x})$ is computed in two steps. First, we construct an encoder $g_{\boldsymbol{\eta}}(\boldsymbol{x})$ similar to the one in (2), which estimates $\boldsymbol{c}$ from $\boldsymbol{x}$
$q^{u}_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$ $=\prod_{j=1}^{D}q^{u}_{\boldsymbol{\eta}}(c_{j}|\boldsymbol{x})=\prod_{j=1}^{D}\text{Ber}(c_{j};g_{j,\boldsymbol{\eta}}(\boldsymbol{x})),$ (10)
@@
where the superscript $u$ indicates that this posterior does not utilize the ECC (uncoded). Now, we leverage the known redundancy introduced by the ECC to constrain the solution of $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x})$, given that each bit from $\boldsymbol{m}$ has been repeated $L$ times to create $\boldsymbol{c}$. To do so, we follow a soft decoding approach, where the marginal posteriors of the information bits are derived from the marginal posteriors of the encoded bits, exploiting the code’s known structure. In the case of repetition codes, we compute the all-are-zero and the all-are-ones products of probabilities of the bits in $\boldsymbol{c}$ that are copies of the same message bit and re-normalize as
$\begin{gathered}q(m_{k}=1|\boldsymbol{x})=\frac{1}{Z}\prod_{j=L(k-1)+1}^{Lk}q^{u}_{\boldsymbol{\eta}}(c_{j}=1|\boldsymbol{x}),\\ q(m_{k}=0|\boldsymbol{x})=\frac{1}{Z}\prod_{j=L(k-1)+1}^{Lk}q^{u}_{\boldsymbol{\eta}}(c_{j}=0|\boldsymbol{x}),\end{gathered}$ (11)
for $k=1,\ldots,M$ and $Z$ is the normalization constant. For the $M=3$, $L=2$ toy example in (7)
$q_{\boldsymbol{\eta}}(m_{1}=1|\boldsymbol{x})\propto q_{\boldsymbol{\eta}}^{u}(c_{1}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}^{u}(c_{2}=1|\boldsymbol{x}),$
$q_{\boldsymbol{\eta}}(m_{2}=1|\boldsymbol{x})\propto q_{\boldsymbol{\eta}}^{u}(c_{3}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}^{u}(c_{4}=1|\boldsymbol{x}),$
$q_{\boldsymbol{\eta}}(m_{3}=1|\boldsymbol{x})\propto q_{\boldsymbol{\eta}}^{u}(c_{5}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}^{u}(c_{6}=1|\boldsymbol{x}).$
@@
This approach can be seen as a soft majority voting strategy. All operations in (11) are differentiable and implemented in the log domain. We consider the same architecture for the encoder $g_{\boldsymbol{\eta}}(\boldsymbol{x})$ in (10) as in the uncoded case, with the only difference being the final MLP layer. The additional overhead in the coded case requires $(L-1)\times h^{\prime}$ parameters in the last layer, where $h^{\prime}$ is the output dimension. In Appendix K.2, we conduct an ablation study on the number of trainable parameters to show that performance gains do not stem from the increased parameter count.
@@
Figure 5: Reconstruction performance in FMNIST. The figure at the left shows an example of reconstructed test images obtained with different model configurations. Observe that more details are visualized as we increase the number of bits in the latent space and decrease the coding rate. The table at the right includes reconstruction metrics. Acc is the semantic accuracy, and Conf. Acc the confident semantic accuracy. Entropy is the average entropy of $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ in the test set.
4.2 Efficient reparameterization
Given the variational family in (9), the ELBO remains unchanged from (3). However, the reparameterization trick in (5) assumes independent bits, which does not hold for $\boldsymbol{c}$. To address this issue efficiently during training, we adopt a soft encoding approach. We compute a marginal probability for each bit in $\boldsymbol{c}$, leveraging the ECC structure and the marginal posterior probabilities of the information bits in $\boldsymbol{m}$. For a repetition code, this involves simply replicating the posterior probabilities $q_{\boldsymbol{\eta}}(m_{k}|\boldsymbol{x}),k=1,\ldots,M$, for each copy of the same information bit [^2] For the example in (7), this results in $p_{\boldsymbol{\theta}}(\boldsymbol{c}|\boldsymbol{x})\approx q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})=\text{Ber}(c_{1};q_{1})\text{Ber}(c_{2};q_{1})\text{Ber}(c_{3};q_{2})\text{Ber}(c_{4};q_{2})\text{Ber}(c_{5};q_{3})\text{Ber}(c_{6};q_{3})$..Hence, we treat the bits in $\boldsymbol{c}$ as independent but distributed according to $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. The training algorithm can be found in Appendix B.
When marginalizing $\boldsymbol{c}$ using the soft encoding marginals, we disregard potential correlations between the coded bits. Consequently, sampling from the marginals may produce an invalid codeword. However, since we do not sample the coded bits during training but instead propagate their marginal probabilities, we consider this method to have minimal negative impact. In fact, it can be seen as a form of probabilistic dropout, which enhances robustness during training. Since encoded words may contain inconsistent bits, the decoder learns to utilize the correlated inputs from repeated bits rather than disregarding them. Remark that when sampling from the generative model in test time, we use hard bits encoded into valid codewords, yielding visually appealing samples, indicating effective training.
5 Results
This section empirically evaluates the Coded-DVAE with repetition codes and its uncoded counterpart on reconstruction and generation tasks. In particular, we present results for MNIST [Deng, 2012], FMNIST [Xiao et al., 2017], CIFAR10 [Krizhevsky et al., 2009], and Tiny ImageNet [Le and Yang, 2015]. Additionally, we compare the coded model to the uncoded DVAE trained with the IWAE objective (see Fig. 1 and Appendix H). All experiments use a fixed architecture (see Appendix C), modifying the encoder’s output and decoder’s input to accommodate the repetition code. We identify the models based on the number of information bits for uncoded models and the code rate for coded models.
@@ 5.1 Generation
We first evaluate the models using the image generation task. Fig. 4 presents examples of randomly generated images using various model configurations on FMNIST. Additional results for other datasets are provided in the Supplementary Material. All models produce more detailed and diverse images as the number of information bits increases. However, when the number of latent vectors becomes too high for the dataset’s complexity, some codebook entries remain unspecialized during training. This results in generation artifacts, where images contain overlapping features from different object classes. A visual inspection of Fig. 4 suggests that such artifacts occur more frequently in the uncoded model.
@@
Error metrics. The effectiveness of the ECC-based inference can be assessed by generating images and measuring errors in recovering $\boldsymbol{m}$ by either drawing one sample from the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ or using the Maximum a Posteriori (MAP) estimate. As shown in the table included in Fig. 4, the coded models significantly reduce both Bit Error Rate (BER) and Word Error Rate (WER) compared to the uncoded case for the same number of latent bits. Note also that the error rates grow with the number of latent bits, which is expected due to the increased complexity of the inference process.
Log-likelihood (LL). Aditionally, we estimated the data LL using importance sampling with 300 samples per observation. As shown in the table included in Fig. 4, coded models consistently outperform their uncoded counterparts on both train and test sets, aligning with the rest of the results. LL values generally improve with more information bits, reflecting increased model flexibility. However, reducing the code rate does not enhance LL, suggesting possible decoder overfitting, where reconstruction improves, but LL declines. This may be due to the use of feed-forward networks at the decoder’s input, which might not fully capture the correlations in coded words (see Appendix L for details).
5.2 Reconstruction
We also assess the model’s performance in image reconstruction. In the table in Fig. 5, we report the Peak Signal-To-Noise Ratio (PSNR) of reconstructions on the FMNIST test set, calculated as
$\text{PSNR}=20\log\Big{(}\frac{\text{max}(x_{i,j})}{\text{RMSE}(\boldsymbol{x},\boldsymbol{x}^{\prime})}\Big{)},$ (12)
where $\text{max}(x_{i,j})$ is the highest pixel value, RMSE represents the Root Mean Square Error, and $\boldsymbol{x}^{\prime}$ the reconstructed image. Results for the other datasets can be found in the Supplementary Material. In all cases, coded models outperform their uncoded counterparts in reconstruction quality, as indicated by higher PSNR values, a trend also visible in Fig. 5. As the number of information bits increases, PSNR improves, reflecting greater model flexibility to capture data structure. We also observe improved PSNR with reduced code rates, as more redundancy is added. However, this increased redundancy does not enhance the model’s flexibility, which is determined by the number of information bits.
Figure 6: Example of erroneous reconstructions in FMNIST. The first column in each image shows the original images, while the second column displays the reconstructions. The $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ probability is indicated in each row.
Semantic Accuracy. Since the PSNR operates at the pixel level, it does not account for the semantic errors committed by the model. To address this, we evaluate reconstruction accuracy in Fig. 5, checking whether the model reconstructs images within the correct semantic class (type of clothing in FMNIST). For this, we trained an image classifier for each dataset and compared the predicted labels of the reconstructed images with the ground truth labels. We also report a confident reconstruction accuracy, which is calculated by considering only the images projected into a latent vector with posterior probability above 0.4  [^3] We do not count errors when the MAP of $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})<0.4$.. Results indicate that incorporating an ECC enables the model to produce latent spaces that better capture the semantics of the data.
Posterior Uncertainty Calibration.   Finally, we report the average entropy of the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ over the test set in the table included in Fig. 5. The low entropy in the uncoded models indicates low uncertainty when the model maps data points into the latent space, which could be beneficial if the model consistently assigns high probability to the correct latent vectors. However, the semantic accuracy results show that this is not the case for the uncoded model, as it often projects images into vectors corresponding to the wrong semantic class with high confidence.
Figure 7: Reconstruction results for CIFAR10 (top) and Tiny ImageNet (bottom).